Grade 12th passDiscuss with colleagues and IITians

A heater boils certain amount of water in 15 min.another heater boils same amount of water in 10 min.time taken to boil same amount of water when both are used in parallel is

satya mohanty
8 Years agoGrade 12th pass

1 Answer

Arun
6 Years ago
Given that First heater boils certain amount of water in 15 minutes.
 
Another heater boils it in 10 minutes.
 
Let water boiled by both of them in one minute be 1 / x
 
Amount boiled by first heater in 1 minute = 1 / 15
 
Water boiled by another heater = 1 / 10
 
so the equation forms in this way :
 
1 / 15 + 1 / 10 = 1 / x
 
L.C.M = 30
 
1 × 2 / 15 × 2 + 3 × 1 / 10 × 3 = 1 / x
 
2 + 3 / 30 = 1 / x
 
5 / 30 = 1 / x
 
Cross multiply the fractions
 
5 × x = 1 × 30
 
5x = 30
 
x = 30 / 5
 
x = 6
 
So time taken by both the heaters to boil the water is 6 minutes.
 
So the answer is 6 minutes.